OpenStudy (anonymous):

Two electric bulbs whose resistances are in the ratio of 1 : 2 are connected in parallel to a constant voltage source. The powers dissipated in them have the ratio. a 1 : 2 b 1 : 1 c 2 : 1 d 1 : 4

OpenStudy (anonymous):

P = V^2/R So, the smaller the resistance, the higher the power. Since you had 1 to 2 resistances, you now will have 2 to 1 power dissipations

OpenStudy (anonymous):

to be more clear about why it's 2:1, basically, imagine that P_1 is dissipated by R_1. To find P_2, simply replace R_2 by R_1 in the equation : \[P_2 = V^2/R_2 = V^2/(2R_1) = 0.5 (V^2/R_1) = 0.5 P_1\]

OpenStudy (anonymous):